Find the volume of the solid obtained by rotating the region enclosed by  𝑥=√6sin(𝑦)

 and 𝑥=0

 about the 𝑦-

axis over the interval 0≤𝑦≤𝜋.



1
Expert's answer
2022-05-03T16:31:40-0400


Volume can be calculated by the formula:

"V=\pi\int_0^\pi(\sqrt{6\sin{y}})^2dy=\pi\int_0^\pi6\sin{y}dy=""-6\pi\cos{y}|_0^\pi=-6\pi(-1-1)=12\pi"

Answer: "V=12\pi" .